OpenStudy (anonymous):

\[(x + \frac{1}{x})^2 - \frac{3}{2}(x - \frac{1}{x}) = 4\]

OpenStudy (anonymous):

expand it first multiply both sides by a multiple of x to get rid of the denominator move everything to one side then attempt to factor

OpenStudy (anonymous):

Finally you will get : \[(x - \frac{1}{x})^2 - \frac{3}{2}(x - \frac{1}{x}) = 0\]

OpenStudy (anonymous):

\[x - \frac{1}{x} = 0 \implies x^2 = 1\]

OpenStudy (anonymous):

Or : \[x - \frac{1}{x} = \frac{3}{2}\]

OpenStudy (anonymous):

@waterineyes whered the 4 go?

OpenStudy (anonymous):

\[(x + \frac{1}{x})^2 - 4 = (x - \frac{1}{x})^2\]

OpenStudy (anonymous):

oh wait is a simple substitution \[u=x+\frac{1}{x}\] from there its a quadratic equation solve for u then solve for x

OpenStudy (anonymous):

Then what will you put for x - 1/x ??